OpenStudy (anonymous):

Solve for "K" in the following polynomials given the binomial that is a factor of the polynomial (using synthetic division) x^3 - kx^2 + 54x -20 (x-5)

OpenStudy (accessdenied):

5 | 1 -k 54 -20 5 -5k + 25 -25k + 395 --------------------------------------------- 1 -k + 5 -5k + 79 -25k + 375 dividing the original polynomial by (x-5) should yield a 0-remainder. -25k + 375 = 0 -25k = -375 k = 15

OpenStudy (anonymous):

thank you so much!